Threshold Corporation uses three components in the manufacture of each of its finished products. The three components are called A, B, and C. The master production schedule for the finished product is given below:

Week

1

2

3

4

5

6

7

8

Gross Requirement

100

70

140

85

60

160

110

75

Two units of Component B. are used in each of the finished product and 350 units of B are available on-hand at the beginning of the plan. Given an average weekly demand of 200 units, an order cost of $100, and a holding cost of $1 per unit-week, complete the above table using the EOQ lot sizing rule. Start developing an MRP for car. Note that there are no previously scheduled receipts. Lead-time for Component B is two weeks. Also compute the total increment cost (setup + holding) for the 8-week schedule.
Answer #1

Ordering cost (A) = $100

Weekly demand (D) = 200 units

Holding cost (h) = $1 per unit week

EOQ = sqrt (2*-A*D/h) = sqrt(2*100*200/1) = 200 units

Holding cost for EOQ = EOQ/2 * h = 200/2*1 = $100

Ordering cost at EOQ = (Weekly demand / EOQ) * Ordering cost = (200/200)*100 = $100

Total incremental cost at EOQ = Holding cost for EOQ + Ordering cost at EOQ = 100+100 = $200

1 2 3 4 5 6 7 8
Demand of final product 100 70 140 85 60 160 110 75
1 unit of final product requires 2 unit of B
Requirement B from final product 200 140 280 170 120 320 220 150

we will use EOQ of 200 units when we order.

Opening Inventory (350) 0 1 2 3 4 5 6 7 8 9 10 11
Demand of B 200 140 280 170 120 320 220 150
Scheduled Receipts (0)
Net Inventory before Production 350 350 10 0 0 40 0 0 0
Planned Receipts 400 200 0 400 200 200
Projected Inventory 150 10 130 160 40 120 100 150
MPS order 400 200 400 200 200

Since we ordering 1400 units, which is equal to 7 EOQ's,

therefore the total incremental cost of ordering during 8 week period = 7 * Total incremental cost at EOQ

= 7*200 =$1400